Tau on sol B/5111 was **** . Sunset tomorrow is about 16:41:44 (HLST). Updated sol B/5112 15:27:56 HLST (Mon, 11 Jun 2018 19:13:38 UTC). Switch rovers.
Note that sols 1237, 5107, and 5109 are based on solar array results, not Pancam opacities. Pancam opacities had a lower limit of n/a, 3.5, and 7, respectively. Sol 5111 data is based on Pancam R2 filter (750 nm) thumbnail.
Disclaimer: Information on this page is based on quick-look data products available at the time the page was created. Results may change as new products become available, even many sols after the original processing, and will change as the calibration is validated for archiving. All information is approximate. In case of urgency, normal operational deliveries are faster than this site. And, yes, I know the MY can show up out of order, I am just ignoring that right now.
This page is (or is not, depending on the date above) maintained by Dr. Mark Lemmon of the Mars Exploration Rover science team and the Department of Atmospheric Sciences, College of Geosciences, Texas A&M University.